OpenStudy (anonymous):

simplify: 4/y^2-36 - 9/y+6 divided by 8/y^2-36 - 5/y-6

OpenStudy (callisto):

4/y^2-36 - 9/y+6 = 4/(y-6)(y+6) - 9/y+6 = 4/(y-6)(y+6)- 9(y-6)/(y+6)(y-6) = (-9y+58)/(y-6)(y+6) 8/y^2-36 - 5/y-6 = 8/(y-6)(y+6) - 5(y+6)/(y-6)(y+6) = (-5y-22)/(y-6)(y+6) therefore 4/y^2-36 - 9/y+6 divided by 8/y^2-36 - 5/y-6 = [ (-9y+58)/(y-6)(y+6)] / (-5y-22)/(y-6)(y+6) = (-9y+58) / (-5y-22) = (9y-58)/(5y+22)

OpenStudy (callisto):

i might have made some mistakes here, it's too complicated

OpenStudy (anonymous):

I know! driving me nuts! Ill let you know.....

OpenStudy (anonymous):

You were on point! Thank you so much for your help this am!

OpenStudy (callisto):

welcome